OpenStudy (anonymous):

A company is expanding its building area from 18,000 square feet to 24,660 square feet. What is the percentage increase of the area of the building space?

OpenStudy (anonymous):

no like what percent is it.

OpenStudy (anonymous):

like this 90%

OpenStudy (anonymous):

convert your answer to percent.

OpenStudy (anonymous):

(24660-18000)x100 = \[\frac{ 24660-18000 }{ 18000 } \times100\]= 37%

OpenStudy (anonymous):

ignore my first line

OpenStudy (anonymous):

Thank you very much.

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

the question will close now.